# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ 9 Preparatory Results
$\delta_{1}[a_{11},\ldots,a_{1k_{1}}]=\sigma(a_{11},\ldots,a_{1k_{1}})$,   $\sigma(a_{11},\ldots,a_{1k_{1}})\dot{\rightarrow}c\in R\langle{E\cup F},\emptyset\rangle$,   and   $\sigma(a_{11},\ldots,a_{1k_{1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}c$.
[…]
$\sigma(b_{11},\ldots,b_{1k_{1}})\dot{\rightarrow}b\in R\langle E,F\rangle$, where $b\subseteq c$.
Since $\delta_{1}[b_{11},\ldots,b_{1k_{1}}]=\sigma(b_{11},\ldots,b_{1k_{1}})$, we have
$\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}b$, where $b\in C\langle E,F\rangle$ and $b\subseteq c$.
@@
    $\delta_{n+1}[a_{(n+1)1},\ldots,a_{(n+1)k_{n+1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[a_{(n-1)1},\ldots,a_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[a_{11},\ldots,a_{1k_{0}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}c$,
[…]
@@
[…]
@@
    $\delta_{n+1}[a_{(n+1)1},\ldots,a_{(n+1)k_{n+1}}]=\delta_{n}[a_{(n+1)1},\ldots,a_{(n+1)(j-1)},\sigma(a_{(n+1)j},\ldots,a_{(n+1)(j+m-1)}),a_{(n+1)(j+m)},\ldots,a_{(n+1)k_{n+1}}]$.
@@
    For every $\ell\in\{1,\ldots,j-1\,\}$, $a_{n\ell}=a_{(n+1)\ell}$ and for every $\ell\in\{j+1,\ldots,k_{n}\,\}$, $a_{n\ell}=a_{(n+1)(\ell+m-1)}$.
@@
    In the last reduction step of (a’) we apply the rule $\sigma(a_{(n+1)j},\ldots,a_{(n+1)(j+m-1)})\dot{\rightarrow}a_{nj}\in R\langle{E\cup F},\emptyset\rangle$.
By Condition ($\mathrm{a}^{\prime}$), $R\langle E\cup F,\emptyset\rangle$ reach $c$ from a proper extension of $a_{nj}$. By Condition ($\mathrm{b}^{\prime}$), for every $\ell\in\{\,j,\ldots,j+m-1\,\}$, $b_{(n+1)\ell}\in C\langle E,F\rangle$ and $b_{(n+1)\ell}\subseteq a_{n\ell}$. We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $c\in C\langle E\cup F,\emptyset\rangle$. Therefore by the rewrite rule appearing in Condition (f) there exists a rewrite rule
$\sigma(b_{(n+1)j},\ldots,b_{(n+1)(j+m-1)})\dot{\rightarrow}b$ in $R\langle E,F\rangle$ with $b\subseteq a_{nj}$.
For every $\ell\in\{1,\ldots,j-1\,\}$, let $b_{n\ell}=b_{(n+1)\ell}$,
let $b_{nj}=b$,
for every $\ell\in\{j+1,\ldots,k_{n}\,\}$, let $b_{n\ell}=b_{(n+1)(\ell+j-1)}$.
@@
for every { $\ell\in\{\,1,\ldots,k_{n}\,\}$, $b_{n\ell}\subseteq a_{n\ell}$ and $\delta_{n+1}[b_{(n+1)1},\ldots,b_{(n+1)k_{n+1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n}[b_{n1},\ldots,b_{nk_{n}}]$.
By the sequence of reductions $(a^{\prime})$ we have
$\delta_{n}[a_{n1},\ldots,a_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[a_{(n-1)1},\ldots,a_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[a_{21},\ldots,a_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{1}[a_{11},\ldots,a_{1k_{0}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}c$.
@@
[…]
there exists $d\in C\langle E,F\rangle$ with $d\subseteq c$
@@
$\delta_{n}[b_{n1},\ldots,b_{nk_{n}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-1}[b_{(n-1)1},\ldots,b_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\cdots\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}d$.
@@
$\delta_{n+1}[b_{(n+1)1},\ldots,b_{(n+1)k_{n+1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n}[b_{n1},\ldots,b_{nk_{n}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-1}[b_{(n-1)1},\ldots,b_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\newline \cdots\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{1}[b_{11},\ldots,b_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}d$.∎
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant $c\in C\langle E\cup F,\emptyset\rangle$. Let $s\in T_{\Sigma}$ and $s\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}c$. Then there exists $d\in C\langle E,F\rangle$ with $d\subseteq c$ such that $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}d$.
Proof.
Lemma 10.97 implies the lemma. ∎
@@
()($\Rightarrow$)( ⇒ ) Let $t\in W$. Then by our assumption,
$t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle E,F\rangle}\mbox{ or }t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle F,E\rangle}$.
@@
$t_{1}/_{\Theta\langle E,F\rangle}=b_{1},\ldots,t_{m}/_{\Theta\langle E,F\rangle}=b_{m}$,
$t_{1}/_{\Theta\langle{E\cup F},\emptyset\rangle}=a_{1},\ldots,t_{m}/_{\Theta\langle{E\cup F},\emptyset\rangle}=a_{m}$.
Hence $\sigma(t_{1}/_{\Theta\langle{E\cup F},\emptyset\rangle},\ldots,t_{m}/_{\Theta\langle{E\cup F},\emptyset\rangle})\dot{\rightarrow}c\in R\langle{E\cup F},\emptyset\rangle$.
@@
$\delta[\sigma(t_{1},\ldots,t_{m})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\delta[\sigma(a_{1},\ldots,a_{m})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta[c]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
@@
$\delta[\sigma(t_{1},\ldots,t_{m})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t_{\Theta\langle E,F\rangle}$.
By Proposition 4.31, $R\langle E,F\rangle$ is convergent. Observe that $t_{\Theta\langle E,F\rangle}$ is irreducible for $R\langle E,F\rangle$. Thence we have
@@
[…]
@@
()($\Leftarrow$)( ⇐ ) Let $t\in W$. We assumed that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. By Definition 10.91 GTRS $R\langle E,F\rangle$ or GTRS $R\langle F,E\rangle$ simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\in C\langle E\cup F,\emptyset\rangle$. Assume that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$. Then $t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle E,F\rangle}$. Let $s\in t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$. Consequently by Proposition 4.36, $s\mathop{\rightarrow}_{R\langle E\cup F,E\rangle}^{*}t/_{\Theta\langle{E\cup F},\emptyset\rangle}$. By Lemma 10.98, there exists $d\in C\langle E,F\rangle$ such that $d\subseteq t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}d$. Then $d=p/_{\Theta\langle E,F\rangle}$ for some $p\in W$. By Proposition 4.34, $s\mathop{\leftrightarrow}^{*}_{E\cup F}t$ and $s\mathop{\leftrightarrow}^{*}_{E}p$. Then $s\mathop{\leftrightarrow}^{*}_{E\cup F}p$. Thus $d=p/_{\Theta\langle E,F\rangle}\subseteq t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle E,F\rangle}$. Consequently $d=t/_{\Theta\langle E,F\rangle}$. Hence $s\in t/_{\mathop{\leftrightarrow}^{*}_{E}}$. Consequently, for any $s\in t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$, we have $s\in t/_{\mathop{\leftrightarrow}^{*}_{E}}$. Thus $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\subseteq t/_{\mathop{\leftrightarrow}^{*}_{E}}$, and hence $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$. We showed that if GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$, then $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$.
@@
We continue Example 3.21 and its sequels, Examples 5.54, 7.73. There does not exist a path of positive length in $AUX[E;F]$. There exists one cycle in $AUX[E;F]$. The cycle is of length $1$, it starts from the vertex #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.70, $R\langle E\cup F,\emptyset\rangle$ reach #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$ and #$\#/_{\Theta\langle F,E\rangle}.keeps_{F}=\mathrm{false}$. Then the formula
#$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$ or #$\#/_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$
does not hold. Thence the following statement does not hold: for each $a\in C\langle E\cup F,\emptyset\rangle$,
  • $a.equal_{E}=\mathrm{true}$ or $a.equal_{F}=\mathrm{true}$ and
  • if $a\not\in C\langle E,F\rangle$, then for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.

### v2

@@ 9 Preparatory Results
$\delta_{1}[{b}_{11},\ldots,{b}_{1k_{1}}]=\sigma({b}_{11},\ldots,{b}_{1k_{1}})$,   $\sigma({b}_{11},\ldots,{b}_{1k_{1}})\dot{\rightarrow}{d}\in R\langle{E\cup F},\emptyset\rangle$,   and   $\sigma({b}_{11},\ldots,{b}_{1k_{1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{d}$.
By Condition (b), for every $j\in\{\,1,\ldots,k_{1}\,\}$, ${c}_{1j}\in C\langle E,F\rangle$ and ${c}_{1j}\subseteq{b}_{1j}$. We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant ${d}\in C\langle E\cup F,\emptyset\rangle$. Consequently by the rewrite rule $\sigma({b}_{11},\ldots,{b}_{1k_{1}})\dot{\rightarrow}{d}\in R\langle{E\cup F},\emptyset\rangle$ and Definition 10.98, and Statement 10.99, there exists a rewrite rule
$\sigma({c}_{11},\ldots,{c}_{1k_{1}})\dot{\rightarrow}{c}\in R\langle E,F\rangle$, where ${c}\subseteq{d}$.
Since $\delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]=\sigma({c}_{11},\ldots,{c}_{1k_{1}})$, we have
$\delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}{c}$, where ${c}\in C\langle E,F\rangle$ and ${c}\subseteq{d}$.
@@
    $\delta_{n+1}[{b}_{(n+1)1},\ldots,{b}_{(n+1)k_{n+1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n}[{b}_{n1},\ldots,{b}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[{b}_{(n-1)1},\ldots,{b}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[{b}_{21},\ldots,{b}_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{1}[{b}_{11},\ldots,{b}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{d}$,
    where for every $i\in\{\,0,\ldots,n+1\,\}$, $\delta_{i}\in CON_{\Sigma,{k_{i}}}$, $k_{i}\in{\mathbb{N}}$, and ${b}_{i1},\ldots,{b}_{i{k_{i}}}\in C\langle{E\cup F},\emptyset\rangle$.
@@
    We obtain $\delta_{n+1}$ from $\delta_{n}$ by replacing the $j$th occurrence from left to right of the hole $\diamondsuit$ by the context $\sigma(\diamondsuit,\ldots,\diamondsuit)$, where $j\in\{\,1,\ldots,k_{n}\,\}$, $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, and $k_{n}=k_{n+1}-m+1$.
@@
    $\delta_{n+1}[{b}_{(n+1)1},\ldots,{b}_{(n+1)k_{n+1}}]=\delta_{n}[{b}_{(n+1)1},\ldots,{b}_{(n+1)(j-1)},\sigma({b}_{(n+1)j},\ldots,{b}_{(n+1)(j+m-1)}),{b}_{(n+1)(j+m)},\ldots,{b}_{(n+1)k_{n+1}}]$.
@@
    For every $\ell\in\{1,\ldots,j-1\,\}$, ${b}_{n\ell}={b}_{(n+1)\ell}$ and for every $\ell\in\{j+1,\ldots,k_{n}\,\}$, ${b}_{n\ell}={b}_{(n+1)(\ell+m-1)}$.
@@
    In the last reduction step of (a’) we apply the rule $\sigma({b}_{(n+1)j},\ldots,{b}_{(n+1)(j+m-1)})\dot{\rightarrow}{b}_{nj}\in R\langle{E\cup F},\emptyset\rangle$.
By Condition ($\mathrm{a}^{\prime}$), $R\langle E\cup F,\emptyset\rangle$ reaches d{d} from a proper extension of ${b}_{nj}$. By Condition ($\mathrm{b}^{\prime}$), for every $\ell\in\{\,j,\ldots,j+m-1\,\}$, ${c}_{(n+1)\ell}\in C\langle E,F\rangle$ and ${c}_{(n+1)\ell}\subseteq{b}_{n\ell}$. We assumed that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant ${d}\in C\langle E\cup F,\emptyset\rangle$. Therefore by the rewrite rule appearing in Condition (f) there exists a rewrite rule
$\sigma({c}_{(n+1)j},\ldots,{c}_{(n+1)(j+m-1)})\dot{\rightarrow}{c}$ in $R\langle E,F\rangle$ with ${c}\subseteq{b}_{nj}$.
For every $\ell\in\{1,\ldots,j-1\,\}$, let ${c}_{n\ell}={c}_{(n+1)\ell}$,
let ${c}_{nj}={c}$,
for every $\ell\in\{j+1,\ldots,k_{n}\,\}$, let ${c}_{n\ell}={c}_{(n+1)(\ell+j-1)}$.
@@
for every { $\ell\in\{\,1,\ldots,k_{n}\,\}$, ${c}_{n\ell}\subseteq{b}_{n\ell}$ and $\delta_{n+1}[{c}_{(n+1)1},\ldots,{c}_{(n+1)k_{n+1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n}[{c}_{n1},\ldots,{c}_{nk_{n}}]$.
By the sequence of reductions ($\mathrm{a}^{\prime}$) we have
$\delta_{n}[{b}_{n1},\ldots,{b}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{n-1}[{b}_{(n-1)1},\ldots,{b}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\cdots\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta_{2}[{b}_{21},\ldots,{b}_{2k_{2}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\newline \delta_{1}[{b}_{11},\ldots,{b}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{d}$.
@@
for any $i\in\{\,1,\ldots,n-1\,\}$ and { $\ell\in\{\,1,\ldots,k_{i}\,\}$, there exist ${c}_{i\ell}\in C\langle E,F\rangle$, with ${c}_{i\ell}\subseteq{b}_{i\ell}$, and
there exists ${e}\in C\langle E,F\rangle$ with ${e}\subseteq{d}$
@@
$\delta_{n}[{c}_{n1},\ldots,{c}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-1}[{c}_{(n-1)1},\ldots,{c}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\cdots\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}{e}$.
@@
$\delta_{n+1}[{c}_{(n+1)1},\ldots,{c}_{(n+1)k_{n+1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n}[{c}_{n1},\ldots,{c}_{nk_{n}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{n-1}[{c}_{(n-1)1},\ldots,{c}_{(n-1)k_{n-1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}\newline \cdots\mathop{\rightarrow}_{R\langle E,F\rangle}\delta_{1}[{c}_{11},\ldots,{c}_{1k_{1}}]\mathop{\rightarrow}_{R\langle E,F\rangle}{e}$.∎
Lemma 10.107 implies the following lemma.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$ such that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant ${d}\in C\langle E\cup F,\emptyset\rangle$. Let $s\in T_{\Sigma}$ and $s\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}{d}$. Then there exists ${e}\in C\langle E,F\rangle$ with ${e}\subseteq{d}$ such that $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}{e}$.
@@
$(\Rightarrow)$ Let $t\in W$. Then by our assumption,
@@
Assume that $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{{\mathop{\leftrightarrow}^{*}_{E}}}$. The case $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{{\mathop{\leftrightarrow}^{*}_{F}}}$ is similar. We now show that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on $t/_{\Theta\langle E\cup F,\emptyset\rangle}$. To this end, let ${d}\in C\langle E\cup F,\emptyset\rangle$ be such that $R\langle{E\cup F},\emptyset\rangle$ reaches $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ from a proper extension of the constant d{d}. Then there exists $\delta\in CON_{\Sigma,1}$ such that $\delta[{d}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle{E\cup F},\emptyset\rangle}$. Let $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{d}\in R\langle{E\cup F},\emptyset\rangle$ be a rewrite rule with ${b}_{1},\ldots,{b}_{m}\in C\langle{E\cup F},\emptyset\rangle$, and let ${c}_{1},\ldots,{c}_{m}\in C\langle E,F\rangle$ with ${c}_{i}\subseteq{b}_{i}$ for all $i\in\{\,1,\ldots,m\,\}$. We now show that there exists a rewrite rule in $R\langle E,F\rangle$ with left-hand side $\sigma({c}_{1},\ldots,{c}_{m})$. Let $t_{1}\in{c}_{1},\ldots,t_{m}\in{c}_{m}$. By Definition 4.27, we have
@@
$t_{1}/_{\Theta\langle E,F\rangle}={c}_{1},\ldots,t_{m}/_{\Theta\langle E,F\rangle}={c}_{m}$,
$t_{1}/_{\Theta\langle{E\cup F},\emptyset\rangle}={b}_{1},\ldots,t_{m}/_{\Theta\langle{E\cup F},\emptyset\rangle}={b}_{m}$.
Hence $\sigma(t_{1}/_{\Theta\langle{E\cup F},\emptyset\rangle},\ldots,t_{m}/_{\Theta\langle{E\cup F},\emptyset\rangle})\dot{\rightarrow}{d}\in R\langle{E\cup F},\emptyset\rangle$.
@@
$\delta[\sigma(t_{1},\ldots,t_{m})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\delta[\sigma({b}_{1},\ldots,{b}_{m})]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta[{d}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
@@
$\delta[\sigma(t_{1},\ldots,t_{m})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
By Proposition 4.32, $R\langle E,F\rangle$ is convergent. Observe that $t/_{\Theta\langle E,F\rangle}$ is irreducible for $R\langle E,F\rangle$. Thence we have
@@
By the set memberships $t_{1},\ldots,t_{m}\in W$ and by Proposition 4.33, we have $t_{j}\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t_{j}/_{\Theta\langle{E},F\rangle}$ for every $j\in\{\,1,\ldots,m\,\}$. Then
@@
$(\Leftarrow)$ Let $t\in W$. We assumed that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. By Definition 10.101 GTRS $R\langle E,F\rangle$ or GTRS $R\langle F,E\rangle$ simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ on the constant $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\in C\langle E\cup F,\emptyset\rangle$. Assume that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$. Then $t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle E,F\rangle}$. Let $s\in t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$. Consequently by Proposition 4.38, $s\mathop{\rightarrow}_{R\langle E\cup F,E\rangle}^{*}t/_{\Theta\langle{E\cup F},\emptyset\rangle}$. By Lemma 10.108, there exists ${e}\in C\langle E,F\rangle$ such that ${e}\subseteq t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}{e}$. Then ${e}=p/_{\Theta\langle E,F\rangle}$ for some $p\in W$. By Proposition 4.35, $s\mathop{\leftrightarrow}^{*}_{E\cup F}t$ and $s\mathop{\leftrightarrow}^{*}_{E}p$. Then $s\mathop{\leftrightarrow}^{*}_{E\cup F}p$. Thus ${e}=p/_{\Theta\langle E,F\rangle}\subseteq t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle E,F\rangle}$. Consequently ${e}=t/_{\Theta\langle E,F\rangle}$. Hence $s\in t/_{\mathop{\leftrightarrow}^{*}_{E}}$. Consequently, for any $s\in t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$, we have $s\in t/_{\mathop{\leftrightarrow}^{*}_{E}}$. Thus $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\subseteq t/_{\mathop{\leftrightarrow}^{*}_{E}}$, and hence $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$. We showed that if GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$, then $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$.
@@
We continue Example 3.22 and its sequels, Examples 5.55, 7.74. There does not exist a path of positive length in $AUX[E;F]$. There exists one cycle in $AUX[E;F]$. The cycle is of length $1$, it starts from the vertex $_{\Theta\langle E\cup F,\emptyset\rangle}$, and leads to $_{\Theta\langle E\cup F,\emptyset\rangle}$. By Lemma 6.71, $R\langle E\cup F,\emptyset\rangle$ reaches $_{\Theta\langle E\cup F,\emptyset\rangle}$ from a proper extension of $_{\Theta\langle E\cup F,\emptyset\rangle}$. Recall that $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{false}$ and $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{false}$ and $_{\Theta\langle F,E\rangle}.keeps_{F}=\mathrm{false}$. Then the formula
$_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{E}=\mathrm{true}$ or $_{\Theta\langle E\cup F,\emptyset\rangle}.equal_{F}=\mathrm{true}$
does not hold. Thence the following statement does not hold: for each ${b}\in C\langle E\cup F,\emptyset\rangle$,
  • 1.
    ${b}.equal_{E}=\mathrm{true}$ or ${b}.equal_{F}=\mathrm{true}$ and
  • 2.
    if ${b}\not\in C\langle E,F\rangle$, then for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{F}=\mathrm{true}$.